OpenStudy (milo123):

Medal: A standard showerhead in Jen's house dispenses 5 gallons of water per minute. Jen changed this showerhead to an energy-saving one. The equation shows the amount of water dispensed, y, as a function of the number of minutes, x, for the new showerhead. y = 3x How much water does Jen save each day with the change in showerheads if she uses the shower for 8 minutes each day? 16 gallons 17 gallons 32 gallons 37 gallons

OpenStudy (milo123):

@Will.H

OpenStudy (milo123):

@563blackghost

OpenStudy (will.h):

So the original is y = 5x The saving- process system is y = 3x solve for both when x = 8 y = 5x when x = 8 y = 5* 8 y = 40 The other system y = 3x when x = 8 y = 3 * 8 y = 24 The saving is : 40 - 24 = 16gallons

563blackghost (563blackghost):

You first need to determine the amount of water she `used` with the oringinal shower head. We now it uses `5 gallons a minute` so we would multiply that by 8. \(\huge\bf{o=5(8)}\) Once that is found we find the amount used with the new shower head. \(\huge\bf{y=3(8)}\) We would then subtract the two to see how much she saves. \(\huge\bf{o-y=?}\)

563blackghost (563blackghost):

or as @Will.H said ^.^

563blackghost (563blackghost):

Nicely done :)

OpenStudy (will.h):

Thanks ghosty you too :)

563blackghost (563blackghost):

thank you will :)

OpenStudy (milo123):

thanks!!